# A Wide-Sense Stationarity Test Based on the Geometric Structure of Covariance

### v1

@@ Abstract
This paper presents a test for wide-sense stationarity (WSS) based on the geometry of the covariance function. We estimate local patches of the covariance surface and then check whether the directional derivative in the $(1,1,0)$ direction is zero on each patch. The method only requires the covariance function to be locally smooth and does not assume stationarity in advance. It can be used for general stochastic dynamical systems and gives a time-resolved view. We apply the test method to a SDOF system and to a stochastic Duffing oscillator. These examples show that the method is numerically stable and can detect departures from WSS in practice.
@@ 1 Introduction
Wide-sense stationarity (WSS) is a commonly used assumption in time series and stochastic process studies, and it serves as the basis for many frequency domain methods and statistical inference (brockwell2009time; liang2015random). In engineering signals, random vibration, and stochastic dynamical systems, the response is often governed by nonlinear dynamics and driven by non-stationary external excitation, and system parameters may also vary over time. Many frequency domain tools rely on WSS. Therefore, to justify frequency-domain analysis, it is necessary to evaluate whether the process is WSS and, if not, to locate the time intervals over which a WSS approximation remains reasonable(lin2004probabilistic; sun2006stochastic; forgoston2018primer).
@@
For non-stationary processes, existing work follows two lines. The first is the locally stationary model, which treats non-stationarity as a slow time variation and builds tests and measures based on this idea(priestley1965evolutionary; dahlhaus1997fitting; paparoditis2009testing; dette2011measure; nason2013test; aue2020testing; van2021nonparametric). The second is the piecewise stationary model, which views non-stationarity as changes at a few time points or as piecewise changes, and develops a range of tests and estimation methods(page1954continuous; hinkley1971inference; andrews1993tests; yao1987approximating; bai1998estimating; einmahl2003empirical); it can also be extended to handle cases where slow variation and sudden shifts occur together(last2008detecting; casini2024change). A shared problem is that one usually needs to decide whether the main source of non-stationarity is slow variation or sudden shifts. If this choice is wrong, it can introduce bias and may even lead to misleading conclusions (casini2024change). This issue is more common in stochastic dynamical systems: Noise can cause sudden switches between metastable states, so the structure may change abruptly rather than slowly forgoston2018primer.On the other hand, putting an SDE directly into a change-point framework often requires a more explicit model setup, so it may be less reliable for complex systems.
@@
This work builds on a simple fact from differential geometry do2016differential: if a stochastic process is WSS, its covariance surface has a cylindrical structure, with generator direction $(1,1,0)$. We therefore assess WSS by testing the local change of the covariance surface along this direction. Using local polynomial regression fan2018local; masry1996multivariate; de2013derivative, we estimate the required derivatives and test whether the directional derivative along $(1,1,0)$ is zero. The proposed method does not rely on local spectra or a specific parametric time series model. It only requires the covariance surface to be continuous and differentiable. When model information is limited or the system is complex, this geometric approach provides a direct and easy to interpret WSS test. The framework is especially suitable for stochastic differential equations and stochastic dynamical systems. On the one hand, under common conditions, solutions of SDEs have well-behaved properties. On the other hand, it is common to obtain multiple independent sample paths through numerical simulation or repeated experiments which helps stabilize covariance estimation and support local diagnostics.
@@
Section 2 describes the geometric structure of the covariance function for WSS processes and introduces local cylindrification as the key step for constructing the WSS test. Section 3 builds the test statistic via local polynomial regression and establishes its properties. Section 4 evaluates the method on two stochastic dynamical SDOF oscillator and a stochastic Duffing oscillator and further compares its performance with the stationarity measure of dette2011measure.
@@ 2.1 Geometric Shape
Consider the surface $M=\left\{x=(s,t,r(s,t))|s,t\in T\right\}$ which is simply the graph of $r(s,t)$. Based on classical results in differential geometry do2016differential, we have the following lemma. Here we assume $r(s,t)\in C^{2}(T\times T)$ so that the Gaussian curvature is well defined.
@@
If $X_{t}$ is WSS, then $M$ is a cylindrical surfacee and its ruling direction is $(1,1,0)$.
@@
Figure 1 shows such a surface. If the process is WSS, the ruling direction is $(1,1,0)$.The surface is a cylinder. We will use this property to build a test for WSS.
@@ 2.2 Local Cylindrification
Lemma 2.2 shows that there is a direct link between WSS and the geometric structure of the covariance surface. Thus, testing WSS can be turned into checking whether the surface is a cylinder with ruling direction $(1,1,0)$. Interestingly, this detection method can also be applied locally. The following lemma provides a general construction of this approach.
@@
Since $r\in C^{1}(\Omega_{i})$ and $\int_{\Omega_{i}}(r(x)-\bar{r}_{\Omega_{i}})dx=0$, we can apply the result of bebendorf2003note. Let $k=\sup_{x\in T\times T}|\nabla r(x)|$. Then
@@
Note that Lemma 2.3 only establishes the existence of a local cylindrification approximation; it does not specify the ruling direction on each patch. When the original function is a cylindrical surface, the ruling direction produced by this construction coincides with the true ruling direction. However, the original function is not cylindrical and therefore does not possess a ruling in the classical sense. This motivates a stronger construction, built on Lemma 2.4, which both selects an appropriate ’generating direction’ and remains applicable to non-cylindrical surfaces.
@@ 2.3 WSS Testing
Case with first order derivative jumps. In this case, classical partial derivatives cannot be defined on $\Sigma$. We therefore use weak derivatives leoni2017first. Since $r\in W^{1,p}(T\times T)$ with $p>2$, the weak partial derivatives $\frac{\partial r}{\partial s}$ and $\frac{\partial r}{\partial t}$ satisfy for any $\phi\in C_{c}^{\infty}(T\times T)$
@@ 3 Local Polynomial Regression
where $m$ is an unknown smooth function. Local polynomial regression approximates $m$ in a neighbourhood of $\mathbf{x}_{0}$ by a Taylor polynomial of degree $p$
@@
    Good asymptotic behaviour. With suitable bandwidth and sufficiently large sample size, local polynomial regression gives consistent estimators for the regression function and for its derivatives fan2018local. Moreover, the suitably scaled estimation errors are asymptotically normal masry1996multivariate. This is convenient for later statistical inference.
@@ 3.1 Partial derivative estimation
where $E(\epsilon(s,t)_{k}|(s,t))=0$, $Var(\epsilon(s,t)_{k}|(s,t))=\sigma^{2}<\infty$ and $E(\epsilon_{k}^{4})<\infty$. We use the two-dimensional Epanechnikov kernel fan2018local
@@ 3.2 Asymptotic theory
where $n$ is the total number of design points and $h^{2}$ is the area of the window. If the window area is fixed, then the number of sample points depends only on the $n$. Moreover
@@
is a continuous map the continuous mapping theorem billingsley2013convergence gives that $\hat{J}$ is a consistent estimator of $J$. ∎
@@
for some symmetric positive definite matrix $\Sigma$. By a weighted central limit theorem van2000asymptotic
@@
the delta method van2000asymptotic yields
@@ 4.1 SDOF
Although we report the statistic at the diagonal point $(t_{i},t_{i})$ for convenience, it is estimated via a local polynomial fit on the covariance surface using nearby values in $\Omega_{i}$, rather than by pointwise differentiation of the diagonal.Of course, we could introduce small offsets, but under our theoretical framework this is unnecessary and would instead sacrifice computational and inferential convenience.
@@
From Figure 6 we see that the process is clearly non-stationary from zero to about eight seconds. Between about eight seconds and about fifteen seconds, the stationarity result fluctuates. After about sixteen seconds, the process gradually becomes WSS.This is consistent with the known results from stochastic dynamics sun2006stochastic.
If we apply Eq. (8.14) in sun2006stochastic
@@ 4.2 Duffing Oscillator
Consider the Duffing stochastic oscillator belousov2019volterra
@@
Classical results lin2004probabilistic give approximate formulas for the covariance of the response, but here our main interest is how stationarity evolves in time.
@@
For the evolution of the probability density of the Duffing oscillator under Gaussian white noise, there is a rich literature. When $a=0$, there is no stationary probability density belousov2019volterra. This corresponds to Case 2, and the test also reports non-stationarity. Cases 1, 3, and 4 have stationary probability densities, so after a period of time, the responses become stationary. This is in agreement with the result of our test.
We see that Case 3 has two potential wells, while Case 1 and Case 4 each have only one. In the single well case, damping removes energy quickly and the probability density soon collapses to one sharp peak. The response statistics then settle in a short time, and the statistic $\hat{J}(t)$ departs from zero only briefly. The difference between Case 1 and Case 4 comes from the value of a, which changes the rate of convergence PhysRevE.87.062132.
In the double well case (Case 3), the system first stays in one well, so the probability density at early times is unimodal and asymmetric. In the long run the density should be bimodal and almost symmetric. Before frequent transitions between the two wells occur, the process has not yet reached its stationary distribution belousov2019volterra. This leads to a longer transient period in the statistic $\hat{J}(t).$
@@ 4.3 Comparison
We compare the method proposed in dette2011measure with the method proposed in this paper. As shown in Figure 10, both methods can detect the evolution of WSS. The difference is that our method does not rely on the assumption of a locally stationary model.
@@
However, for a Wiener process, the DPV measure is less informative in our setting. As shown in Figure 11, J(t) has the theoretical value 1 and its estimate stays close to 1, clearly indicating a persistent departure from WSS, whereas the DPV measure does not provide a comparably clear diagnostic.
@@ 5 Conclusion
In this paper, we propose a geometric method to test WSS directly. We write the classical WSS condition as a first-order condition on the covariance function. Using a local cylindrification scheme together with local polynomial regression, we build an estimator of the derivatives and show that it is consistent and asymptotically normal. The test does not rely on any fixed parametric time series model. It works directly on the covariance surface and gives geometrically interpretable diagnostics for WSS.
The method also has several limitations. First, the current theory still needs regularity conditions on the covariance function, such as piecewise smoothness and the existence of derivatives up to some finite order. Hence, it does not yet apply to very rough covariance surfaces. Second, since local polynomial regression is performed on the two-dimensional time plane, the algorithm needs a lot of computation and requires large number of observations.
Future research will put more emphasis on spectral representation and data analysis within this framework. For small samples and rough covariance surfaces, it is necessary to develop more robust inference and testing methods, so as to improve finite-sample performance while relaxing smoothness assumptions.For the representation and spectral decomposition of the covariance function, preliminary results show that, if $r$ has continuous first-order partial derivatives, then on a local interval the covariance can be written as:
$r(s,t)=h_{s_{0}}(s-t)+\int_{s_{0}}^{t}J(s-t+\xi,\xi)d\xi,$ (102)
where $J(s,t)=r_{s}(s,t)+r_{t}(s,t)$, $h_{s_{0}}$ describes the corresponding locally wide-sense stationary part, and the integral term captures geometric distortions of non-stationarity along the time direction. On this basis, one can introduce a local spectral density $S_{s_{0}}(\omega,t)$, which leads to a geometrically meaningful evolutionary spectral representation centered around $J(s,t)$. In this paper, we only discuss this representation at a conceptual level; its full theory and the related statistical inference will be developed in future work.

### v2

@@ Abstract
This paper presents a test for wide-sense stationarity based on the geometry of the covariance function. We estimate local patches of the covariance surface and then check whether the directional derivative in the $(1,1,0)$ direction is zero on each patch. The method only requires the covariance function to be locally smooth and does not assume stationarity in advance. It can be applied to general stochastic dynamical systems and provides a time-resolved view. We apply the test method to an SDOF system and to a Duffing oscillator. These examples show that the method is numerically stable and can detect departures from stationarity in practice.
@@ 1 Introduction
Wide-sense stationarity (WSS) is a commonly used assumption in time series and stochastic process studies, and it serves as the basis for many frequency domain methods and statistical inference (Brockwell and Davis (2009); Liang and Lee (2015)). In engineering signals, random vibration, and stochastic dynamical systems, the response is often governed by nonlinear dynamics and driven by non-stationary external excitation, and system parameters may also vary over time. Therefore, to justify frequency domain analysis, it is necessary to evaluate whether the process is WSS and, if not, to locate the time intervals over which a WSS approximation remains reasonable(Lin and Cai (2004); Sun (2006); Forgoston and Moore (2018)).
@@
For non-stationary processes, existing work follows two lines. The first is the locally stationary model, which treats non-stationarity as a slow time variation and builds tests and measures based on this idea(Priestley (1965); Dahlhaus (1997); Paparoditis (2009); Dette et al. (2011); Nason (2013); Aue and Van Delft (2020); van Delft et al. (2021)). The second is the piecewise stationary model, which views non-stationarity as changes at a few time points or as piecewise changes, and develops a range of tests and estimation methods(Page (1954); Hinkley (1971); Andrews (1993); Yao (1987); Bai and Perron (1998); Einmahl and McKeague (2003)); it can also be extended to handle cases where slow variation and sudden shifts occur together(Last and Shumway (2008); Casini and Perron (2024)). A shared problem is that one usually needs to decide whether the main source of non-stationarity is slow variation or sudden shifts. If this choice is wrong, it can introduce bias and may even lead to misleading conclusions (Casini and Perron (2024)). This issue is more common in stochastic dynamical systems: Noise can cause sudden switches between metastable states, so the structure may change abruptly rather than slowly Forgoston and Moore (2018).On the other hand, putting an SDE directly into a change-point framework often requires a more explicit model setup, so it may be less reliable for complex systems.
@@
This work builds on a fact from differential geometry Do Carmo (2016): if a stochastic process is WSS, its covariance surface has a cylindrical structure, with generator direction $(1,1,0)$. We therefore assess WSS by testing the local change of the covariance surface along this direction. Using local polynomial regression Fan (2018); Masry (1996); De Brabanter et al. (2013), we estimate the required derivatives and test whether the directional derivative along $(1,1,0)$ is zero. The proposed method does not rely on local spectra or a specific parametric time series model. It only requires the covariance surface to be continuous and differentiable. When model information is limited, or the system is complex, this geometric approach provides a direct and easy to interpret test. The framework is especially suitable for stochastic differential equations and stochastic dynamical systems. On the one hand, under common conditions, solutions of SDEs have well-behaved properties. On the other hand, it is common to obtain multiple independent sample paths through numerical simulation or repeated experiments, which helps stabilize covariance estimation and support local diagnostics.
@@
Section 2 describes the geometric structure of the covariance function for WSS processes and introduces local cylindrification as the key step for constructing the WSS test. Section 3 builds the test statistic via local polynomial regression and establishes its properties. Section 4 evaluates the method on two stochastic dynamical SDOF oscillators and a stochastic Duffing oscillator and further compares its performance with the stationarity measure of Dette et al. (2011).
@@ 2.1 Geometric Shape
Consider the surface $M=\left\{x=(s,t,r(s,t))|s,t\in T\right\}$ which is simply the graph of $r(s,t)$. Based on classical results in differential geometry Do Carmo (2016), we have the following lemma. Here we assume $r(s,t)\in C^{2}(T\times T)$ so that the Gaussian curvature is well defined.
@@
If $X_{t}$ is WSS, then $M$ is a cylindrical surface, and its ruling direction is $(1,1,0)$.
@@
Figure 1 shows such a surface. If the process is WSS, the ruling direction is $(1,1,0)$. We will use this property to build a test for WSS.
@@ 2.2 Local Cylindrification
Lemma 2.2 shows that there is a direct link between WSS and the geometric structure of the covariance surface. Thus, testing WSS can be turned into checking whether the surface is a cylinder with ruling direction $(1,1,0)$. Interestingly, this detection method can also be applied locally. The following lemma provides a construction of this approach.
@@
Since $r\in C^{1}(\Omega_{i})$ and $\int_{\Omega_{i}}(r(x)-\bar{r}_{\Omega_{i}})dx=0$, we can apply the result of Bebendorf (2003). Let $k=\sup_{x\in T\times T}|\nabla r(x)|$. Then
@@
Note that Lemma 2.3 only establishes the existence of a local cylindrification approximation; it does not specify the ruling direction on each patch. When the original function is a cylindrical surface, the ruling direction produced by this construction with the true ruling direction. However, the original function is not cylindrical and therefore does not possess a ruling in the classical sense. This motivates a stronger construction, built on Lemma 2.4, which both selects an appropriate ’generating direction’ and remains applicable to non-cylindrical surfaces.
@@ 2.3 WSS Testing
Case with first order derivative jumps. In this case, classical partial derivatives cannot be defined on $\Sigma$. We therefore use weak derivatives Leoni (2017). Since $r\in W^{1,p}(T\times T)$ with $p>2$, the weak partial derivatives $\frac{\partial r}{\partial s}$ and $\frac{\partial r}{\partial t}$ satisfy for any $\phi\in C_{c}^{\infty}(T\times T)$
@@ 3 Local Polynomial Regression
where $m$ is an unknown smooth function. Local polynomial regression approximates $m$ in a neighborhood of $\mathbf{x}_{0}$ by a Taylor polynomial of degree $p$
@@
    Good asymptotic behavior. With suitable bandwidth and sufficiently large sample size, local polynomial regression gives consistent estimators for the regression function and for its derivatives Fan (2018). Moreover, the suitably scaled estimation errors are asymptotically normal Masry (1996). This is convenient for later statistical inference.
@@ 3.1 Partial derivative estimation
where $E(\epsilon(s,t)_{k}|(s,t))=0$, $Var(\epsilon(s,t)_{k}|(s,t))=\sigma^{2}<\infty$ and $E(\epsilon_{k}^{4})<\infty$. We use the two-dimensional Epanechnikov kernel Fan (2018)
@@ 3.2 Asymptotic theory
where $n$ is the total number of design points and $h^{2}$ is the area of the window. If the window area is fixed, then the number of sample points depends only on $n$. Moreover
@@
is a continuous map the continuous mapping theorem Billingsley (2013) gives that $\hat{J}$ is a consistent estimator of $J$. ∎
@@
for some symmetric positive definite matrix $\Sigma$. By a weighted central limit theorem Van der Vaart (2000)
@@
the delta method Van der Vaart (2000) yields
@@ 4.1 SDOF
Although we report the statistic at the diagonal point $(t_{i},t_{i})$ for convenience, it is estimated via a local polynomial fit on the covariance surface using nearby values in $\Omega_{i}$, rather than by pointwise differentiation of the diagonal. Of course, we could introduce small offsets, but under our theoretical framework, this is unnecessary and would instead sacrifice computational and inferential convenience.
@@
From Figure 6 we see that the process is clearly non-stationary from zero to about eight seconds. Between about eight seconds and about fifteen seconds, the stationarity result fluctuates. After about sixteen seconds, the process gradually becomes WSS. This is consistent with the known results from stochastic dynamics Sun (2006).
If we apply Eq. (8.14) in Sun (2006)
@@ 4.2 Duffing Oscillator
Consider the Duffing stochastic oscillator Belousov et al. (2019)
@@
Classical results Lin and Cai (2004) give approximate formulas for the covariance of the response, but here our main interest is how stationarity evolves in time.
@@
For the evolution of the probability density of the Duffing oscillator under Gaussian white noise, there is a rich literature. When $a=0$, there is no stationary probability density Belousov et al. (2019). This corresponds to Case 2, and the test also reports non-stationarity. Cases 1, 3, and 4 have stationary probability densities, so after a period of time, the responses become stationary. This is in agreement with the result of our test.
We see that Case 3 has two potential wells, while Case 1 and Case 4 each have only one. In the single well case, damping removes energy quickly, and the probability density soon collapses to one sharp peak. The response statistics then settle in a short time, and the statistic $\hat{J}(t)$ departs from zero only briefly. The difference between Case 1 and Case 4 comes from the value of a, which changes the rate of convergence Méndez et al. (2013).
In the double well case (Case 3), the system first stays in one well, so the probability density at early times is unimodal and asymmetric. In the long run, the density should be bimodal and almost symmetric. Before frequent transitions between the two wells occur, the process has not yet reached its stationary distribution Belousov et al. (2019). This leads to a longer transient period in the statistic $\hat{J}(t).$
@@ 4.3 Comparison
We compare the method proposed in Dette et al. (2011) with the method proposed in this paper. As shown in Figure 10, both methods can detect the evolution of WSS. The difference is that our method does not rely on the assumption of a locally stationary model.
@@
However, for a Wiener process, the DPV measure is less informative in our setting. As shown in Figure 11, J(t) has the theoretical value 1, and its estimate stays close to 1, clearly indicating a persistent departure from WSS, whereas the DPV measure does not provide a comparably clear diagnostic.
@@ 5 Conclusion
In this paper, we propose a geometric method to test WSS directly. We write the classical WSS condition as a first-order condition on the covariance function. Using a local cylindrification together with local polynomial regression, we build an estimator of the derivatives and show that it is consistent and asymptotically normal. The test does not rely on any fixed parametric time series model. It works directly on the covariance surface and gives geometrically interpretable diagnostics for WSS.
The method also has several limitations. First, the current theory still needs regularity conditions on the covariance function, such as piecewise smoothness and the existence of derivatives up to some finite order. Hence, it does not yet apply to very rough covariance surfaces. Second, since local polynomial regression is performed on the two-dimensional time plane, the algorithm needs a lot of computation and requires a large number of observations.
Future research will put more emphasis on spectral representation and data analysis within this framework. For small samples and rough covariance surfaces, it is necessary to develop more robust inference and testing methods to improve finite-sample performance while relaxing smoothness assumptions. For the representation and spectral decomposition of the covariance function, we will consider providing a spectral representation of the covariance function for non-stationary stochastic processes from an evolutionary perspective, enabling better frequency-domain analysis of the responses of stochastic dynamical systems.